### v1

# Small time asymptotics of spectral heat content of isotropic processes

Abstract.
We provide a general approach for proving small time asymptotic of spectral heat content for any translation invariant isotropic process satisfying negligible tail probability condition. As a consequence, we recover several existing results in the context of Lévy processes and Gaussian processes, and provide spectral heat content asymptotic for a class of time-changed Brownian motions.
Key words and phrases:
spectral heat content; isotropic processes; asymptotic behavior
2020 Mathematics Subject Classification:
60J45; 60G51; 60G15
1. Introduction and main result
Let $X=(X_{t})_{t\geqslant 0}$ be a stochastic process on $\mathbb{R}^{d}$ with càdlàg paths and $\mathbb{P}_{x}$ denotes the probability measure when $X_{0}=x$. For any bounded open set $\Omega\subset\mathbb{R}^{d}$, the spectral heat content of $X$ in the domain $\Omega$ is defined as
$Q^{X}_{\Omega}(t)=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}>t)dx,$
where $\tau^{X}_{\Omega}=\inf\{t\geqslant 0:X_{t}\notin\Omega\}$. Since $X$ has càdlàg paths, it follows that
$\lim_{t\to 0}Q^{X}_{\Omega}(t)=\mathrm{Vol}(\Omega),$
where $\mathrm{Vol}(\Omega)$ denotes the volume of $\Omega$. In this article, we consider translation invariant isotropic processes $X=(X_{t})_{t\geqslant 0}$ on $\mathbb{R}^{d}$, that is, for any $x\in\mathbb{R}^{d}$, and $A\in\mathcal{B}(\mathbb{R}^{d})$,
$\mathbb{P}_{x}(X_{t}\in A)$ $=\mathbb{P}_{0}(X_{t}+x\in A),$
$\mathbb{P}_{0}(X_{t}\in A)$ $=\mathbb{P}_{0}(UX_{t}\in A)$
for all $d\times d$ orthogonal matrices $U$. Writing $X_{t}=(X^{(1)}_{t},\ldots,X^{(d)}_{t})$, for any $t>0$, we define
(1.1) $\mu(t)$ $=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge 1\right].$
From right continuity of the sample paths we have
$\lim_{t\to 0}\mu(t)=0.$
Throughout the article, we assume the negligible tail probability condition.
Assumption 1.1.
For any $\varepsilon>0$,
$\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)=0.$
The goal of this article is to prove the following asymptotic of the spectral heat content of $X$.
Theorem 1.1.
Let $\Omega$ be bounded open subset of $\mathbb{R}^{d}$ with $C^{1,1}$-boundary, and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic stochastic process satisfying Assumption 1.1. Then,
(1.2) $\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\mathrm{Per}(\Omega),$
where $\mathrm{Per}(\Omega)$ denotes the perimeter of $\Omega$.
Remark 1.2.
The $C^{1,1}$ regularity condition on the boundary of $\Omega$ inherits from the differentiability of the signed distance function from the boundary, see Remark 6.6 for details.
Corollary 1.3.
Let $\Omega$ be a bounded, open set with $C^{1,1}$-boundary and $X=(X_{t})_{t\geqslant 0}$ be a translation invariant isotropic process such that
$\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant T}X^{(1)}_{s}\right)^{\alpha}\right]<\infty$
for some $T>0$ and $\alpha>1$, and
$\lim_{t\to 0}\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right]}=0.$
Then,
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega),$
where $m(t)=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right$].
When $X=B$, the standard Brownian motion on $\mathbb{R}^{d}$, the spectral heat content can be realized in terms of the Dirichlet heat equation. More precisely, if $u(t,x)$ solves the PDE
$\frac{\partial u}{\partial t}u(x,t)$ $=\frac{1}{2}\Delta u(x,t)\quad\mbox{in $\Omega\times(0,\infty)$}$
$u(x,t)$ $=0\quad\mbox{for $(x,t)\in\partial\Omega\times(0,\infty)$}$
$u(x,0)$ $=1\quad\mbox{for $x\in\Omega$},$
then the spectral heat content on $\Omega$ for $B$ is given by
$Q^{B}_{\Omega}(t)=\int_{\Omega}u(x,t)dx.$
Small time asymptotic of $Q^{B}_{\Omega}(t)$ was first proved by Berg, see [12, Theorem 6.2], for domains with $R$-smooth boundary, and subsequently, a second order expansion of $Q^{B}_{\Omega}(t)$ near zero was proved in [14, Theorem 1.2] for domains with $C^{3}$-boundary. We also refer to [13] for small time asymptotic of $Q^{B}_{\Omega}$ when $B$ is the Brownian motion on a Riemannian manifold and $\Omega$ has smooth boundary. The spectral heat content asymptotic for symmetric $\alpha$-stable was obtained in [9], and it was later generalized to isotropic Lévy processes in [4], and to symmetric Lévy processes in a recent article [6]. We also refer to [5] for the spectral heat content asymptotic of Gaussian processes in $\mathbb{R}^{d}$.
In the present article, we provide a general approach for spectral heat content asymptotic that works for any translation invariant isotropic process satisfying Assumption 1.1. This assumption holds true for any isotropic Lévy process with regularly varying Lévy-Khintchine exponent, and Gaussian processes. Moreover, in Section 2.3, we provide examples of isotropic processes which are neither Markov nor Gaussian and Theorem 1.1 holds for such processes.
Our method also extends to bounded domains with boundary being level sets of a $C^{1,\kappa}$ functions for $0<\kappa<1$. In all the aforementioned references [12, 13, 14, 9, 4, 5], the assumption of $C^{1,1}$-regularity is crucial as the proofs rely heavily on the uniform interior and exterior ball condition of the domain. In fact, from [7, Corollary 2], the uniform interior and exterior ball condition is equivalent to the $C^{1,1}$ smoothness of the boundary. For $C^{1,\kappa}$-boundary with $0<\kappa<1$, the uniform ball condition does not hold. Theorem 6.1 provides upper and lower bound of the spectral heat content for domains with boundaries given by level sets of $C^{1,\kappa}$ functions with $0<\kappa<1$. The question of validity of (1.2) for such domains still remains open.
We propose a novel approach by considering a functional version of spectral heat content. For any nonnegative integrable function $f:\mathbb{R}^{d}\to\mathbb{R}$ we define
(1.3) $Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx,$
In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$ for all $t\geqslant 0$. First, we obtain asymptotic formula for $Q^{X}_{f}$ when $f$ is a compactly supported smooth function, and the lower bound of the limit in (1.2) is proved using an approximation argument. It turns out that the lower bound of (1.2) holds for any bounded, open subset with finite perimeter. The main idea behind proving the upper bound of the limit in (1.2) stems from the observation that any bounded domain with $C^{1,1}$-boundary can be expressed as $\phi^{-1}(0,\infty)$ for some bounded $C^{1,1}$ function $\phi$ with $|\nabla\phi|$ being constant on the boundary.
The rest of the paper is organized as follows: in Section 2 we provide examples including isotropic Lévy processes, Gaussian processes, and a class of time-changed Brownian motions for which spectral heat content asymptotic follow directly from our main result. Section 3 provides some preliminaries about functions of bounded variation and regularity of boundary. In Section 4, we prove the asymptotic of the functional of the spectral heat content defined in (1.3). Finally, Theorem 1.1 is proved in Section 6.
2. Examples
2.1. Isotropic Lévy processes
We consider the class of Lévy processes studied in [4]. Let $X=(X_{t})_{t\geqslant 0}$ be an isotropic Lévy process with Lévy-Khintchine exponent $\Psi(\xi)=\psi(|\xi|)$ for any $\xi\in\mathbb{R}^{d}$, and assume that $\psi:[0,\infty)\to[0,\infty)$ is regularly varying with exponent $\beta\in[1,2]$. Then, by [10, Equation (3.2)], for any $\varepsilon>0$,
$\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)\leqslant c_{\varepsilon}t$
for some constant $c_{\varepsilon}>0$. On the other hand, using the bounds in [6, Equation (3.8)], we have $\lim_{t\to 0}\frac{t}{\mu(t)}=0$, which shows that Assumption 1.1 holds. Therefore, Theorem 1.1 holds for $X$, and we recover [6, Theorem 2.1(1)] for isotropic Lévy processes. In particular, when $X_{t}$ is the symmetric $\alpha$-stable process, as $t\to 0$ we have
$\mu(t)\sim\begin{cases}t^{1/\alpha}\mathbb{E}[\sup_{0\leqslant s\leqslant 1}X^{(1)}_{s}]&\mbox{if $1<\alpha\leqslant 2$}\\ \frac{1}{\pi}t\log(1/t)&\mbox{if $\alpha=1$}.\end{cases}$
For $\alpha\in(1,2]$, the asymptotic behavior of $\mu(t)$ above follows directly from the self-similarity of $\alpha$-stable process, and for $\alpha=1$, we refer to [2, Proposition 4.3(i)]. This recovers [9, Theorem 1.1] when $\alpha\geqslant 1$.
2.2. Gaussian processes
Consider a translation invariant Gaussian process $X=(X_{t})_{t\geqslant 0}$ on $\mathbb{R}^{d}$ with independent and identically distributed components and assume that $X$ has càdlàg sample paths. For any $t>0$, let
$m(t)=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right],\quad\sigma^{2}(t)=\sup_{0\leqslant s\leqslant t}\mathrm{Var}\left[(X^{(1)}_{s})^{2}\right].$
It is known that $m(t)<\infty$ for all $t\geqslant 0$. By Borell-TIS inequality, see [3, Theorem 2.1.1], for any $x>0$
(2.1) $\mathbb{P}_{0}\left(\left|\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}-m(t)\right|>x\right)$ $\leqslant 2e^{-x^{2}/2\sigma^{2}(t)},$
and hence
$\mathrm{Var}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right]\leqslant\int_{0}^{\infty}4xe^{-x^{2}/2\sigma^{2}(t)}=4\sigma^{2}(t).$
Therefore, $\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{2}\right]\leqslant m(t)^{2}+4\sigma^{2}(t)$. From [4, Equation (29)] it is known that $\lim_{t\to 0}{\sigma^{2}(t)}/{m(t)}=0$. Applying Corollary 1.3, we obtain the following result.
Proposition 2.1.
If $\Omega$ is a bounded open set with $C^{1,1}$-boundary then,
(2.2) $\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega).$
Remark $2.$
In [4, Theorem 4.4] under the condition that the rescaled process $Y^{(t)}_{u}=\mu^{-1}_{t}X^{(1)}_{tu}$ converges weakly to a Gaussian process $(Y_{u})_{u\geqslant 0}$ as $t\downarrow 0$ in Skorokhod $J_{1}$ topology, it is proved that
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathbb{E}\left[\sup_{0\leqslant u\leqslant 1}Y_{u}\right]\mathrm{Per}(\Omega),$
with $\mathbb{E}[\sup_{0\leqslant u\leqslant 1}Y_{u}]\leqslant 1$. Proposition 2.1 strengthens [4, Theorem 4.4].
2.3. Time changed Brownian motion
Consider the $d$-dimensional standard Brownian motion $B=(B_{t})_{t\geqslant 0}$ and a nonnegative, increasing continuous process $U=(U_{t})_{t\geqslant 0}$ which is independent of $B$ and $U_{0}=0$. Define $X_{t}=B_{U_{t}}$. Then, $X=(X_{t})_{t\geqslant 0}$ is translation invariant and isotropic. Let us assume that for some $\alpha>1/2$,
(2.3) $\mathbb{E}[U^{\alpha}_{t}]<\infty\quad\forall t\geqslant 0,\quad\lim_{t\to 0}\frac{\mathbb{E}[U^{\alpha}_{t}]}{\mathbb{E}[\sqrt{U_{t}}]}=0.$
Then we have the following result.
Proposition 2.3.
For any bounded open set $\Omega$ with $C^{1,1}$-boundary we have
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}[\sqrt{U_{t}}]}=\sqrt{\frac{2}{\pi}}\mathrm{Per}(\Omega).$
Proof.
Since $U_{t}$ is increasing and continuous, for any $t\geqslant 0$ we have
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}B^{(1)}_{U_{t}}\right]=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant U_{t}}B^{(1)}_{s}\right]=\int_{0}^{\infty}\mathbb{E}_{0}\left[\sup_{0\leqslant u\leqslant s}B^{(1)}_{u}\right]\mathbb{P}(U_{t}\in ds).$
Recalling that for any $p>0$,
$\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}B^{(1)}_{s}\right)^{p}\right]=\frac{2^{p/2}t^{p/2}\Gamma(\frac{p+1}{2})}{\sqrt{\pi}}$
we get
(2.4) $\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}B^{(1)}_{U_{s}}\right)^{p}\right]=\frac{2^{p/2}\Gamma(\frac{p+1}{2})}{\sqrt{\pi}}\mathbb{E}[U^{p/2}_{t}].$
Therefore, (2.3) implies that the conditions of Corollary 1.3 hold. This concludes the proof. ∎
As an example of $U_{t}$, let us consider an $n$-dimensional standard Brownian motion $W=(W_{t})_{t\geqslant 0}$ independent of $B$ and define the integrated squared radial process
$R_{t}=\int_{0}^{t}|W_{s}|^{2}ds.$
Then, the one-dimensional marginal distributions of $X_{t}=B_{R_{t}}$ coincide with the Lévy stochastic area, see [11, Equation (3.5)]. Also, $X$ is neither Markov nor Gaussian. For any $t\geqslant 0$,
$\mathbb{E}[R_{t}]=\int_{0}^{t}nsds=\frac{nt^{2}}{2},$
and by Cauchy-Schwarz inequality $\mathbb{E}[\sqrt{R_{t}}]\leqslant\sqrt{\mathbb{E}[R_{t}]}<\infty$. On the other hand, using Cauchy-Schwarz inequality, we have
$\sqrt{\frac{R_{t}}{t}}\geqslant\frac{1}{t}\int_{0}^{t}|W_{s}|ds,$
which shows that
$\mathbb{E}[\sqrt{R_{t}}]\geqslant\frac{1}{\sqrt{t}}\mathbb{E}\left[\int_{0}^{t}|W_{s}|ds\right]\geqslant ct$
for some constant $c>0$. Therefore, $\lim_{t\to 0}\mathbb{E}[R_{t}]/\mathbb{E}[\sqrt{R_{t}}]=0$, and hence the conditions in (2.3) are satisfied by $R_{t}$, which implies the following corollary.
Corollary 2.4.
Let the processes $B,R$ be defined as above and $X_{t}=B_{R_{t}}$. Then, for any bounded, open set $\Omega$ with $C^{1,1}$-boundary
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}[\sqrt{R_{t}}]}=\sqrt{\frac{2}{\pi}}\mathrm{Per}(\Omega).$
3. Preliminaries: Perimeter and regularity of the boundary
A function $f\in L^{1}(\mathbb{R}^{d})$ has bounded variation if
$\|Df\|:=\sup\left\{\int_{\mathbb{R}^{d}}f\sum_{i=1}^{d}D_{i}\varphi_{i}:\varphi_{i}\in C^{\infty}_{c}(\mathbb{R}^{d}),\sum_{i=1}^{d}\varphi^{2}_{i}\leqslant 1\right\}<\infty.$
In this case, $\|Df\|$ is called the variation of $f$, and the space of all functions with bounded variation is denoted by $\mathrm{BV}(\mathbb{R}^{d})$. A bounded open set $\Omega\subset\mathbb{R}^{d}$ is said to have finite perimeter if $\mathbbm{1}_{\Omega}\in\mathrm{BV}(\mathbb{R}^{d})$, and we write
$\mathrm{Per}(\Omega)=\|D\mathbbm{1}_{\Omega}\|.$
Next, for $\kappa>0$, a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{1,\kappa}$ if $f$ is $C^{1}$ and $\nabla f$ is uniformly Hölder continuous with exponent $\kappa$, that is, there exists $L>0$ such that
$|\nabla f(x)-\nabla f(y)|\leqslant L|x-y|^{\kappa}$
for all $x,y\in\mathbb{R}^{d}$. The domain $\Omega\subset\mathbb{R}^{d}$ has $C^{1,\kappa}$-boundary if for every $x\in\partial\Omega$, there exists a neighborhood $V_{x}$ and a $C^{1,\kappa}$ function $\gamma:\mathbb{R}^{d-1}\to\mathbb{R}$ such that
$\Omega\cap V_{x}=\{(y_{1},\ldots,y_{d}):y_{d}>\gamma(y_{1},\ldots,y_{d-1})\}\cap V_{x}.$
If $\Omega$ has $C^{1}$-boundary, from Stoke’s theorem it follows that $\mathrm{Per}(\Omega)=\mathcal{H}^{d-1}(\partial\Omega)$, the (d1)(d-1)-dimensional Hausdorff measure of $\partial\Omega$.
4. The minimum functional
Consider a nonnegative measurable function $f$ on $\mathbb{R}^{d}$. For any $u\geqslant 0$, let us define
$E^{f}_{u}=\{x:f(x)>u\}.$
We consider the following functional version of the spectral heat content: for any $f$ as above, we define
$Q^{X}_{f}(t)=\int_{0}^{\infty}Q^{X}_{E^{f}_{u}}(t)du=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\mathbb{P}_{x}(\tau^{X}_{E^{f}_{u}}>t)dxdu.$
Lemma 4.1.
For any nonnegative measurable function $f$ on $\mathbb{R}^{d}$ one has
$Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$.
Proof.
We note that for any fixed $u\geqslant 0$,
$\left\{u<\inf_{0\leqslant s\leqslant t}f(X_{s})\right\}\subseteq\{\tau_{E^{f}_{u}}>t\}=\left\{\inf_{s\geqslant 0}\{s:f(X_{s})\leqslant u\}>t\right\}\subseteq\left\{u\leqslant\inf_{0\leqslant s\leqslant t}f(X_{s})\right\}$
Applying Fubini’s theorem,
$Q^{X}_{f}(t)=\int_{\mathbb{R}^{d}}\int_{0}^{\infty}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}f(X_{s})\geqslant u\right)dudx=\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
This completes the proof of the lemma. ∎
In what follows, we will study the asymptotic behavior of the quantity
$\mathcal{R}^{X}_{f}(t):=\int_{\mathbb{R}^{d}}f(x)dx-Q^{X}_{f}(t).$
as $t\to 0$. Trivially, when $f=\mathbbm{1}_{\Omega}$, $\mathcal{R}^{X}_{f}(t)=\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)$.
Theorem 4.2.
Let $X=(X_{t})_{t\geqslant 0}$ be a translation invariant, isotropic stochastic process on $\mathbb{R}^{d}$ satisfying Assumption 1.1. Then for any nonnegative $f\in C^{1,\kappa}_{c}(\mathbb{R}^{d})$,
$\lim_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
where $\mu(t)$ is defined in (1.1). In particular, the above limit holds for all compactly supported smooth functions.
We require the following lemmas to prove the above theorem.
Lemma 4.3.
For any $\varepsilon>0$,
(4.1) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]=1$
(4.2) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\right\}\right]=1$
(4.3) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]=1$
(4.4) $\limsup_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]\leqslant 2d$
Proof.
For any $\varepsilon>0$ we note that
$\left|\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]-\mu(t)\right|\leqslant\max\{\varepsilon,1\}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\wedge 1\right).$
Therefore, (4.1) follows as $\lim_{t\to 0}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\wedge 1\right)/\mu(t)=0$. Next, for any $t>0$, we have
$0$ $\leqslant\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]-\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\right\}\right]$
$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right),$
and hence, (4.2) follows from (4.1) and Assumption 1.1. To prove (4.3), we observe that
$0$ $\leqslant\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\}\right]-\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
$=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon,\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\}\right]$
$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\right)\leqslant 2d\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon/\sqrt{d}\right),$
where the last inequality follows from the isotropy of $X_{t}$. Again by Assumption 1.1 and (4.2), we conclude that
(4.5) $\ \ \ \ \lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
$=\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\}\right]=1.$
Finally, using the symmetry of the random variables $X^{(i)}_{t}$ for $i=1,\ldots,d$, we note that
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\leqslant\varepsilon\}\right]\leqslant\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|>u\right)du$
$\leqslant 2\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(i)}_{s}>u\right)=2\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(i)}_{s}\wedge\varepsilon\right].$
[…]
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]\leqslant\sum_{i=1}^{d}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X^{(i)}_{s}|\leqslant\varepsilon\}\right].$
Therefore, (4.4) follows from (4.5) and (4.3). ∎
Lemma 4.4.
Suppose that $f\in C^{1,\kappa}(\mathbb{R}^{d})$ with $|\nabla f(x)-\nabla f(y)|\leqslant L|x-y|^{\kappa}$ for all $x,y\in\mathbb{R}^{d}$, and $t>0$. Then,
$\left|\inf_{0\leqslant s\leqslant t}f(X_{s})-f(x)-\inf_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla f(x)\rangle\right|\leqslant\frac{L}{1+\kappa}\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}$
Proof.
Since $f\in C^{1}(\mathbb{R}^{d})$, using Taylor expansion around $x$, we obtain,
(4.6) $f(X_{s})=f(x)+\langle X_{s}-x,\nabla f(x)\rangle+\omega_{f}(x,X_{s})$
where $\omega_{f}(x,x^{\prime})=\int_{0}^{1}\langle(x^{\prime}-x),(\nabla f(x+u(x^{\prime}-x))-\nabla f(x))\rangle du$. Using the $\kappa$-Hölder continuity of $\nabla f$, it follows that
(4.7) $|\omega_{f}(x,x^{\prime})|\leqslant\frac{L}{1+\kappa}|x-x^{\prime}|^{1+\kappa}.$
This completes the proof of the lemma. ∎
Proof of Theorem 4.2.
Let us write $S_{f}=\mathrm{supp}(f)$. Then we observe that
$f(x)-\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]=0$
whenever $x\notin S_{f}$. As a result, using (4.6), $\mathcal{R}^{X}_{f}(t)$ reduces to
(4.8) $\mathcal{R}^{X}_{f}(t)$ $=\int_{S_{f}}\mathbb{E}_{x}\left[-\inf_{0\leqslant s\leqslant t}\left(\langle X_{s}-x,\nabla f(x)\rangle+\omega_{f}(X_{s},x)\right)\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx$
Assuming that $S_{f}\subset B_{R}(0)$ for some $R>0$, (4.8) implies
(4.9) $\left|\mathcal{R}^{X}_{f}(t)-\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx\right|$
$\leqslant\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|\omega_{f}(X_{s},x)|\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx$
$\leqslant\frac{L}{1+\kappa}\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx,$
where we used the bound in (4.7). For $\varepsilon>0$, the translation invariance and isotropy of $X$ leads to
$\ \ \ \ \mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]$
$\leqslant\varepsilon^{\kappa}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]+2^{1+\kappa}R^{1+\kappa}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\right).$
This shows with Assumption 1.1 and (4.4) that for any $\varepsilon>0$,
$\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx\leqslant 2d\varepsilon^{\kappa}\mathrm{Vol}(S_{f}).$
Letting $\varepsilon\to 0$, the above inequality combined with (4.9) yields
(4.10) $\lim_{t\to 0}\frac{1}{\mu(t)}\left|\mathcal{R}^{X}_{f}(t)-\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx\right|=0.$
Let us denote
$\delta(x)=\inf\{|x-y|:y\in\partial S_{f}\}.$
Then, any $x\in S_{f}$ and $t\geqslant 0$ we have
(4.11) $\ \ \ \sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant\delta(x)\}$
$\leqslant\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}$
$\leqslant\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\}$
Recalling the isotropy and translation invariance of $X$, for any $r>0$ we have
$\ \ \ \ \mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant r\}\right]$
$=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant r\}\right]|\nabla f(x)|=:\mu(t,r)|\nabla f(x)|.$
As a result, (4.11) implies the following upper and lower bound:
(4.12) $\mu(t,\delta(x))|\nabla f(x)|$
$\leqslant\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s},x\in S_{f},0\leqslant s\leqslant t\}\right]$
$\leqslant\mu(t,2R)|\nabla f(x)|.$
Combining (4.10) and (4.12), we obtain
(4.13) $\limsup_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}$ $\leqslant\limsup_{t\to 0}\frac{\mu(t,2R)}{\mu(t)}\int_{\Omega}|\nabla f(x)|dx$
$=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
where the last equality follows from (4.2) in Lemma 4.3. On the other hand, using (4.10) and (4.12) once again we get
$\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\liminf_{t\to 0}\int_{S_{f}}|\nabla f(x)|\frac{\mu(t,\delta(x))}{\mu(t)}dx.$
Since by (4.2) in Lemma 4.3, $\frac{\mu(t,\delta(x))}{\mu(t)}\to 1$ as $t\to 0$, applying Fatou’s lemma we conclude that
(4.14) $\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\int_{S_{f}}|\nabla f(x)|dx=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx.$
Combining (4.13) and (4.14), we conclude the proof of the theorem. ∎
5. Proof of the lower bound
Proposition 5.1.
Let $\Omega$ be a bounded, open set in $\mathbb{R}^{d}$ with finite perimeter. Then,
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\mathrm{Per}(\Omega).$
Remark 5.2.
When $\Omega$ has $C^{1,1}$-boundary, the above proposition can be proved using a half-space argument similar to the proof of [9, Lemma 3.1]. However, the argument in the aforementioned paper fails when $\partial\Omega$ is not regular.
The above result is proved using approximation argument using mollifiers. Fix a collection of mollifiers $(\rho_{\varepsilon})_{\varepsilon>0}\in C^{\infty}_{c}(\mathbb{R}^{d})$ such that $\int_{\mathbb{R}^{d}}\rho_{\varepsilon}(x)dx=1$ and for any $\varepsilon>0$ and $f\in L^{1}(\mathbb{R}^{d})$, define the convolution
$f_{\varepsilon}=f\ast\rho_{\varepsilon}.$
Lemma 5.3.
For any $f\in L^{1}(\mathbb{R}^{d})$ with $f\geqslant 0$ and $\varepsilon>0$, we have
$\mathcal{R}^{X}_{f_{\varepsilon}}(t)\leqslant\mathcal{R}^{X}_{f}(t)$
for all $t\geqslant 0$, where $f_{\varepsilon}=f\ast\rho_{\varepsilon}$.
Proof.
We note that $\int_{\mathbb{R}^{d}}f_{\varepsilon}(x)dx=\int_{\mathbb{R}^{d}}f(x)dx$ for all $\varepsilon>0$. Also, by translation invariance of $X_{t}$ and Fubini’s theorem we get
$\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f_{\varepsilon}(X_{s})\right]$ $=\mathbb{E}_{0}\left[\inf_{0\leqslant s\leqslant t}f_{\varepsilon}(X_{s}+x)\right]$
$=\mathbb{E}_{0}\left[\inf_{0\leqslant s\leqslant t}\int_{\mathbb{R}^{d}}f(X_{s}+x-y)\rho_{\varepsilon}(y)dy\right]$
$\geqslant\mathbb{E}_{0}\left[\int_{\mathbb{R}^{d}}\inf_{0\leqslant s\leqslant t}f(X_{s}+x-y)\rho_{\varepsilon}(y)dy\right]$
$=\int_{\mathbb{R}^{d}}\mathbb{E}_{x-y}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]\rho_{\varepsilon}(y)dy.$
This shows that for all $\varepsilon>0$,
$\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f_{\varepsilon}(X_{s})\right]dx\geqslant\int_{\mathbb{R}^{d}}\mathbb{E}_{x}\left[\inf_{0\leqslant s\leqslant t}f(X_{s})\right]dx.$
This completes the proof of the lemma. ∎
Proof of Proposition 5.1.
Let us consider $f=\mathbbm{1}_{\Omega}$. By Lemma 5.3 and Theorem 4.2, for any $\varepsilon>0$,
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f_{\varepsilon}}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|dx.$
As $\lim_{\varepsilon\to 0}\|f_{\varepsilon}-\mathbbm{1}_{\Omega}\|_{L^{1}}=0$, by lower semi-continuity of the variation it follows that
$\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\geqslant\liminf_{\varepsilon\to 0}\int_{\mathbb{R}^{d}}|\nabla f_{\varepsilon}(x)|\mu(dx)\geqslant\|D\mathbbm{1}_{\Omega}\|=\mathrm{Per}(\Omega).$
This completes the proof of the proposition. ∎
6. Proof of Theorem 1.1
We prove a stronger version of Theorem 1.1.
Theorem 6.1.
Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a $C^{1,\kappa}$ function with $0<\kappa<1$ and assume that
  1. (i)
    $\Omega=\phi^{-1}(0,\infty)$ is pre-compact,
  2. (ii)
    $|\nabla\phi|$ is bounded away from 0 in a neighborhood of $\partial\Omega$.
Then for any translation invariant isotropic process $X=(X_{t})_{t\geqslant 0}$ satisfying Assumption 1.1 we have
$\mathrm{Per}(\Omega)\leqslant\liminf_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
The theorem is proved in several intermediate steps. First, we note the following result about continuity of the perimeters of level sets.
Lemma 6.2.
Let $\phi:\mathbb{R}^{d}\to\mathbb{R}$ be a $C^{1}$ function such that
  1. (i)
    $\Omega=\phi^{-1}(0,\infty)$ is pre-compact,
  2. (ii)
    $|\nabla\phi|$ is bounded away from 0 in a neighborhood of $\partial\Omega$.
For any $r>0$, define $\Omega_{r}=\phi^{-1}(r,\infty)$. Then,
$\lim_{r\to 0}\mathcal{H}^{d-1}(\partial\Omega_{r})=\mathcal{H}^{d-1}(\partial\Omega)$
Notation 6.1.
For the rest of the article, we denote
$\Omega^{\prime}_{r}=\{x:0<\phi(x)<r\},\quad\Omega_{r}=\{x:\phi(x)>r\},\quad r>0.$
Then, $\partial\Omega_{r}=\phi^{-1}(r)$.
For a proof of the above lemma, we refer the reader to the Mathoverflow post [1] and answer by Mathoverflow user Kostya_I, where the lemma has been proved for smooth functions $\phi$, but the same argument works for $C^{1}$ functions as the proof only relies on inverse function theorem. For the sake of completeness, we provide the proof below with some minor modifications.
Proof.
We claim that every point $x\in\phi^{-1}(0)$ has a neighborhood $V_{x}$ that is a diffeomorphic image $\varphi_{x}(Q_{r_{x}})$ of a box $Q_{r_{x}}=\{|y_{i}|\leqslant r_{x},\;1\leqslant i\leqslant d\}$, such that $\phi(\varphi(y))=|\nabla\phi(x)|\cdot y_{1}$, and such that $\varphi_{x}(0)=x$ and $D\varphi_{x}(0)$ is a rotation. Indeed, without loss of generality we may assume that $x=0$, and, by rotating, that $\nabla\phi(0)/|\nabla\phi(0)|=e_{1},$ the first basis vector. Consider the function $g:\mathbb{R}^{d}\to\mathbb{R}^{d}$ defined by $g(x)=(\phi(x)/|\nabla\phi(0)|,x_{2},\dots,x_{d})$. Then $Dg(0)$ is the identity, and $\varphi_{x}$ is the inverse $g^{-1}$ provided by the inverse function theorem. Given $\varepsilon>0$, by choosing smaller $r_{x}$ if necessary, we can ensure that the restriction of $\varphi_{x}$ to the leaves $\{y_{1}=r\}$ distorts the (d1)(d-1)-area by no more than $1+\varepsilon$, that is, in $Q_{\delta_{x}}$,
(6.1) $1-\varepsilon<\left(\det_{2\leqslant i,j\leqslant d}\left(\partial_{y_{i}}\varphi\cdot\partial_{y_{j}}\varphi\right)\right)^{\frac{1}{2}}\leqslant 1+\varepsilon.$
By compactness, we can choose a finite cover $V_{i}=V_{x_{i}}$, $i=1,\dots,n$ of $\phi^{-1}(0)$. Then, for $\varepsilon_{0}>0$ small enough, we have $\overline{\Omega}^{\prime}_{\varepsilon_{0}}:=\{x:0\leqslant\phi(x)\leqslant\varepsilon_{0}\}\subset\cup_{i=1}^{n}V_{i}$. Let $V_{1},\dots,V_{n}$ be a finite cover of $\Omega^{\prime}_{\varepsilon_{0}}$, and we can pick a partition of unity $f_{0},\dots,f_{n}$ subordinate to that cover. We have for $0<r<\varepsilon_{0}$,
$\mathcal{H}^{d-1}(\phi^{-1}(r))=\sum_{i=1}^{n}\int_{\phi^{-1}(r)}f_{i}(w)\mathcal{H}^{d-1}(dw).$
Hence, by changing the variable and (6.1),
(6.2) $(1-\varepsilon)\sum_{i=1}^{n}\int_{Q^{r}_{i}}f_{i}\circ\varphi_{i}dy_{2}\ldots dy_{d}\leqslant\mathcal{H}^{d-1}(\phi^{-1}(r))\leqslant(1+\varepsilon)\sum_{i=1}^{n}\int_{Q^{r}_{i}}f_{i}\circ\varphi_{i}dy_{2}\ldots dy_{d},$
where $Q^{r}_{i}=Q_{i}\cap\{y_{1}=r\}$. As $r\to 0$, the terms in the right-hand side tend to
$(1+\varepsilon)\int_{Q^{0}_{i}}f_{i}\circ\varphi_{i}dy_{2}\ldots dy_{d}\leqslant(1+\varepsilon)^{2}\int_{\phi^{-1}(0)}f_{i}(w)\mathcal{H}^{d-1}(dw),$
and similarly for the left hand side of (6.2). This means that
$(1-\varepsilon)^{2}\mathcal{H}^{d-1}(\phi^{-1}(0))$ $\leqslant\liminf_{r\to 0}\mathcal{H}^{d-1}(\phi^{-1}(r)),$
$\limsup_{r\to 0}\mathcal{H}^{d-1}(\phi^{-1}(r))$ $\leqslant(1+\varepsilon)^{2}\mathcal{H}^{d-1}(\phi^{-1}(0)),$
and since $\varepsilon>0$ is arbitrary and $\phi^{-1}(r)=\partial\Omega_{r}$ for any $r\geqslant 0$, the proof the lemma follows. ∎
Lemma 6.3.
For any $\varepsilon>0$,
$\lim_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx=0$
Proof.
Invoking Lemma 4.4 we have
$\inf_{0\leqslant s\leqslant t}\phi(X_{s})\geqslant\phi(x)+\inf_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle-\frac{L}{1+\kappa}\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}.$
Therefore,
$\ \ \ \ \mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)$
$\leqslant\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle+\frac{L}{1+\kappa}\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>\phi(x)\right)$
$\leqslant\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|>\frac{\phi(x)}{2M}\right)+\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|>\left(\frac{\phi(x)}{L^{\prime}}\right)^{1/1+\kappa}\right),$
where $M=\|\nabla\phi\|_{\infty}$ and $L^{\prime}=L/(1+\kappa)$. As a result, for any $x\in\Omega_{\varepsilon}$, Assumption 1.1 implies that
$\ \ \ \ \lim_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$\leqslant\lim_{t\to 0}\frac{1}{\mu(t)}\left(\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\frac{\varepsilon}{2M}\right)+\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\left(\frac{\varepsilon}{L^{\prime}}\right)^{1/1+\kappa}\right)\right)\mathrm{Vol}(\Omega)$
$=0.$
This completes the proof of the lemma. ∎
Lemma 6.4.
For any $\eta\in(0,1)$ and $\varepsilon>0$,
$\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}+2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}},$
where $M_{\varepsilon}=\sup_{x\in\Omega^{\prime}_{\varepsilon}}|\nabla\phi(x)|$ and $m_{\varepsilon}=\inf_{x\in\Omega^{\prime}_{\varepsilon}}|\nabla\phi(x)|$.
Proof.
Fix any $\eta\in(0,1)$. For any $x\in\Omega^{\prime}_{\varepsilon}$, using Lemma 4.4 again we get
$\ \ \ \ \mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)$
(6.3) $\leqslant\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)+\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right).$
Using translation invariance and isotropy of $X$, and the co-area formula, integral of the first term on $\Omega^{\prime}_{\varepsilon}$ reduces to
(6.4) $\ \ \ \ \int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
$=\int_{0}^{\varepsilon}\int_{\partial\Omega_{v}}\mathbb{P}_{0}\left(|\nabla\phi(x)|\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\eta v\right)\frac{\mathcal{H}^{d-1}(dx)}{|\nabla\phi(x)|}dv.$
Then, (6.4) yields
$\ \ \ \ \int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{1}{m_{\varepsilon}}\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(M_{\varepsilon}\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\eta v\right)dv$
$=\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}\mathbb{E}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\frac{\eta\varepsilon}{M_{\varepsilon}}\right)$
By (4.1) in Lemma 4.3, we get
(6.5) $\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}\langle X_{s}-x,\nabla\phi(x)\rangle>\eta\phi(x)\right)dx$
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}$
To estimate the integral of the second term in (6.3), we again use co-area formula as follows:
$\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
$=\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
$=\int_{0}^{\varepsilon}\int_{\partial\Omega_{v}}\frac{\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}>(1-\eta)v\right)}{|\nabla\phi(x)|}\mathcal{H}^{d-1}(dx)dv$
$\leqslant\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{1}{(1-\eta)m_{\varepsilon}}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}\wedge(1-\eta)\varepsilon\right]$
(6.6) $\leqslant 2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\mu(t)\varepsilon^{\frac{\kappa}{\kappa+1}},$
where in the last inequality we used
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|^{1+\kappa}\wedge(1-\eta)\varepsilon\right]=\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right)^{\kappa+1}\right]$
$\leqslant((1-\eta)\varepsilon)^{\frac{1}{\kappa+1}}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\wedge(1-\eta)^{\frac{1}{1+\kappa}}\varepsilon^{\frac{1}{1+\kappa}}\right]\leqslant 2\mu(t)((1-\eta)\varepsilon)^{\frac{\kappa}{\kappa+1}}$
for any $\varepsilon\in(0,1)$. Since $m_{\varepsilon}$ is bounded away from 0 as $\varepsilon\to 0$, (6.6) implies
(6.7) $\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}>(1-\eta)\phi(x)\right)dx$
$\leqslant 2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}}.$
Hence, the proof of the lemma follows after combining (6.5) and (6.7). ∎
Proof of Theorem 6.1.
By Proposition 5.1, it suffices to prove that
$\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
For any $\varepsilon>0$ we have
(6.8) $\ \ \ \ \limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}=\int_{\Omega}\mathbb{P}_{x}(\tau^{X}_{\Omega}\leqslant t)dx$
$\leqslant\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx$
$+\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{\Omega^{\prime}_{\varepsilon}}\mathbb{P}_{x}\left(\inf_{0\leqslant s\leqslant t}\phi(X_{s})\leqslant 0\right)dx.$
By Lemma 6.3, the first term on the right hand side of (6.8) is zero. Recalling Lemma 6.4, we get
$\ \ \ \ \limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}$
$\leqslant 2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}+2\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}}$
As $|\nabla\phi|$ is bounded in $\Omega^{\prime}_{\varepsilon}$ for sufficiently small $\varepsilon>0$, using Lemma 6.2, we have
$\lim_{\varepsilon\to 0}\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{M_{\varepsilon}}{\eta m_{\varepsilon}}=\frac{1}{\eta}\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathcal{H}^{d-1}(\partial\Omega),$
$\lim_{\varepsilon\to 0}\sup_{0\leqslant v\leqslant\varepsilon}\mathcal{H}^{d-1}(\partial\Omega_{v})\frac{(1-\eta)^{-\frac{\kappa}{\kappa+1}}}{m_{\varepsilon}}\varepsilon^{\frac{\kappa}{\kappa+1}}=0.$
Finally, letting $\eta\to 1$, we conclude that
$\limsup_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mu(t)}\leqslant\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathcal{H}^{d-1}(\partial\Omega)=\frac{\sup_{\partial\Omega}|\nabla\phi|}{\inf_{\partial\Omega}|\nabla\phi|}\mathrm{Per}(\Omega).$
This completes the proof of the theorem. ∎
Concluding proof of Theorem 1.1
The final step of the proof of Theorem 1.1 is to construct a function $\phi$ satisfying conditions of Theorem 6.1 for domains with $C^{1,1}$-boundary.
Lemma 6.5.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set with $C^{1,1}$-boundary. Then there exists a bounded $C^{1,1}$ function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ such that
  1. (1)
    $\Omega=\{x:\phi(x)>0\}$,
  2. (2)
    $|\nabla\phi(x)|=1$ in a neighborhood of $\partial\Omega$.
Proof.
Consider the signed distance function
$\delta(x)=\begin{cases}d(x,\partial\Omega)&\mbox{if $x\in\Omega$}\\ -d(x,\partial\Omega)&\mbox{if $x\in\Omega^{c}$}.\end{cases}$
Since $\Omega$ has $C^{1,1}$-boundary, there exists $r>0$ such that $\delta$ is $C^{1,1}$ in $\Omega^{\prime}_{r}:=\{x\in\mathbb{R}^{d}:|\delta(x)|<r\}$ and $|\nabla\delta|=1$ in $\Omega^{\prime}_{r}$. Define a $C^{1,1}$ function $\varphi:\mathbb{R}\to\mathbb{R}$ such that $\varphi(x)=x$ when $|x|<r/2$ and equals $r$ when $|x|>r$ and consider $\phi=\varphi\circ\delta$. Clearly, $\phi$ is $C^{1,1}$ and $\phi(x)>0$ if and only if $x\in\Omega$. Also, for any $x\in\Omega^{\prime}_{r/2}$, $|\nabla\phi(x)|=|\varphi^{\prime}(\delta(x))\nabla\delta(x)|=1$. This completes the proof of the lemma. ∎
Remark 6.6.
The $C^{1,1}$-regularity assumption on $\partial\Omega$ cannot be weakened. From [8, Corollary 6] it is known that the differentiability of $\delta$ in a neighborhood of $\partial\Omega$ implies that $\partial\Omega$ is $C^{1,1}$.
We note that $\phi$ in Lemma 6.5 satisfies the conditions of Theorem 6.1 with $\kappa=1$, and $\Omega=\phi^{-1}(0,\infty)$. Also, $|\nabla\phi|=1$ on $\partial\Omega$. Therefore, Theorem 1.1 follows from Theorem 6.1. ∎
Proof of Corollary 1.3.
For any $\varepsilon>0$, using Markov inequality we obtain
$\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right)\leqslant\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{\varepsilon^{\alpha}}.$
Also, from the given conditions and noting that $\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\geqslant 0$ if $X_{0}=0$, we have
$0\leqslant 1-\frac{\mu(t)}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>1\right\}\right]}{m(t)}\leqslant\frac{\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\right)^{\alpha}\right]}{m(t)},$
and the right hand side goes to 0 as $t\to 0$. Hence, $\mu(t)/m(t)\to 1$ as $t\to 0$ and Assumption 1.1 is satisfied. The proof is concluded by applying Theorem 1.1. ∎